Amanda is 5 years older than Peter and Sam is 3 years younger than Peter.If the sum of their ages is 50, find their respective ages.

Answer by Edwin McCravy(20056)

Amanda is 5 years older than Peter
A = P+5

and Sam is 3 years younger than Peter.
S = P-3

If the sum of their ages is 50,
A+P+S = 50

Substitute (P+5) for A, and (P-3) for S in A+P+S = 50:

(P+5)+P+(P-3) = 50

Solve for P, that's Peter's age.

Substitute what you get for P in A = P+5 to get Amanda's age.

Substitute what you get for P in S = P-3 to get Peters's age.

Edwin